Circular No. 5914

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 N. Bartel and M. Bietenholz, York University; M. Rupen, J. Conway,

T. Beasley, R. Sramek, and J. Romney, National Radio Astronomy Observatory; M. Titus, Haystack Observatory; D. Graham, Max-Planck-Institut für Radioastronomie; V. Altunin and D. Jones, Jet Propulsion Laboratory; A. Rius, Instituto de Astronomia y Geodesia, Madrid; T. Venturi and G. Umana, Instituto di Radioastronomia, CNR; R. Francis, M. McCall, M. Richer, and C. Stevenson, York University; K. Weiler, Naval Research Laboratory; S. Van Dyk, Naval Research Laboratory and University of California at Berkeley; N. Panagia, Space Telescope Science Institute; W. Cannon, Institute for Space and Terrestial Science and York University; J. Popelar, Energy, Mines, and Resources, Ottawa; and R. Davis, Nuffield Radio Astronomy Laboratories, report: “Global VLBI observations of SN 1993J during the first three months after the explosion, with 9 to 15 antennas for up to 18 hours, show a supernova consistent with being circularly symmetric and expanding uniformly. A selected set of 30 percent of the data is now correlated and analyzed. Following are epochs, observing frequencies, and the radii of circular uniform disk models used in fits to the visibility data: Apr. 27.2 UT, 22.2 GHz, 96 ± 8 microarcsec; May 17.2, 22.2, 145 ± 14; May 17.2, 8.4, 151 ± 11; June 27.0, 14.9, 241 ± 21; June 27.0, 8.4, 275 ± 9; June 27.0, 4.9, 205 ± 85. A weighted least-squares linear fit gives a zero-point of expansion of Mar. 25 ± 4. With the more precise optically derived shock-breakout date of Mar. 28.0 ± 0.1 (Wheeler et al. 1993, Ap.J. 417, L71), the expansion rate is 2.98 ± 0.08 microarcsec per day. An equivalent power-law fit gives an index of 0.96 ± 0.07, showing no evidence yet for deceleration of the shock front. Combining the angular expansion rate trigonometrically with the optically derived maximum expansion speed of the hydrogen gas, we obtain a value for the distance to M81 of 4.0 ± 0.6 Mpc. All errors are 1σ uncertainties, with statistical and systematic contributions included. We have continued the observations at intervals of six weeks and should be able to obtain a sequence of detailed images of the expanding supernova.”

NOVA CASSIOPEIAE 1993 Further CCD photometry by H. Mikuž, Ljubljana, Slovenia (cf. IAUC

5912): Dec. 28.72 UT, V = 7.00 ± 0.02, B–V = +0.68 ± 0.04.

1993 December 30 Daniel W. E. Green